# Tokens, the oft-overlooked appetizer: Large language models, the distributional hypothesis, and meaning

### v6

@@ 2.1 Exemplar tokenizations
A tokenization algorithm uses the training data it is given to determine what pieces (tokens) will be needed for future output. The popular tokenization design choice of maximizing orthographic efficiency allows as much data as possible to be processed by the LLM, presumably the motivation behind this strategy’s ubiquity. We trained one such tokenizer, a Byte-Pair Encoding (BPE) tokenizer Anthony Moi and Nicolas Patry (2023), with varying vocabulary sizes and training data to explore the details of the process, demonstrating some alignment between linguistic features and information processing by showing the human-like stages the tokens pass through. As a caveat, the training data chosen also determines the tokens; once vocabulary size becomes relatively large compared to the training data, tokens become too specific.
@@
Figure 2: Tokens passing through word-like stages as the vocabulary size parameter increases, using a BPE tokenizer and the beginning of Alice in Wonderland. Syntax tree via  Christodoulopoulos (2023); GPT-4o tokenization via  Duong (2023).
The tokenizer’s vocabulary size hyperparameter sets the total number of tokens that will be known to the model. Due to the way the BPE tokenization algorithm works, as the vocabulary size increases, the maximum length of the tokens also increases (an initial base vocabulary covering the fundamental units, usually characters or bytes, is expanded through a series of pairwise merges). Therefore, in iteratively increasing the vocabulary size, we must292929in the logic of the Intermediate Value Theorem at some point pass through token lengths that are the lengths of human-meaningful units, e.g., word length, but the tokens do not need to actually be those units, e.g., words.
@@
Figure 3: The tokenization of the first sentence from Alice in Wonderland by a BPE tokenizer trained on the wikitext dataset and that sentence, as the vocabulary size increases from 2,000 to 30,000, then additional much larger vocabularies (the figure is intended to be explorable in a PDF viewer). The tokens at vocabulary size of 1,000 were individual characters; there were around 3 times as many tokens in the sentence as at vocabulary size of 2,000, so imagine a tower on the extreme left that is 3 times as tall as the one shown. Likewise, on the extreme right would eventually be the sentence as a single token. We skipped some vocabulary sizes along the x-axis: it is not evenly spaced (intermediate behaviour between the largest vocabulary sizes was omitted for legibility, but can be reasonably extrapolated by the reader). Middle English translation via  OpenL . The short tokens, while combinatorily potent and requiring the model to memorize only a small vocabulary, engender prolific output: the ratio of total tokens per generated text are essentially as high as possible given the source material. On the other hand, the large tokens, while efficient in terms of total tokens per generated text (in the limit, 1:1), are far too specific (the model would need to memorize an essentially infinite vocabulary in order to output general purpose text).
@@ 2.2 Exemplar vocabularies
GPT-4o OpenAI et al. (2024); Shahriar et al. (2024) is a recent, prominent, proprietary flagship model, and we take its vocabulary to be representative of what a good vocabulary currently looks like. We can see in Figure 2 that GPT-4o uses tokens that look similar to the word-like stage (achieved through linguistics-agnostic tokenizing) and like tokens generated based on linguistic structures (morphemes, lemmas, words), consistent with our expectation.
However, relatively little information is available on what tokens look like more generally out in the wild. An exception is (the very fun) work of Zhemchuzhina et al., 2022, which delves into what kinds of tokens occur in a vocabulary, and comes up with the following categories: (1) atom, the smallest discrete element, which is probably coming from a closed set (as in the letters of an alphabet in written language); (2) pragma, which consists of atoms, represents part of an idea, and has multiple meanings (which seemingly would correspond to morphemes); (3) idea, a sequence of pragmas whose meaning is a single specific concept (which they relate to utterances and phrases). Due to the vocabulary size of most models, they point out that most tokens correspond to atoms and pragmas rather than ideas Zhemchuzhina, Filippov, and Yamshchikov (2022).
We expect that function words (e.g., “from”, “this”), generally short, frequent, and considered a fairly small closed class in English Brysbaert et al. (2016), are therefore more likely to be tokenized (i.e., even occur in small vocabularies) compared to the majority of content words. [^33] This seems to be true as shown in Figure 8.
We extend this work by examining a group of in-use LLM (including MLM) vocabularies to see what their tokens actually look like. We specifically seek commonalities between human language and tokens, since we suspect emerging best practices should be nudging vocabularies in the human-like direction, but our primary goal is to gain insight into what the tokens in real, reasonably widely-used models look like.
@@ 2.3 Extispicy: exemplar tokens
Previous work found semantic information was a dominant force within the GPT-3 LLM’s internal organization of language. Digutsch and Kosinski, 2023, found that “GPT-3’s semantic activation is better predicted by similarity in words’ meaning (i.e., semantic similarity) rather than their co-occurrence in the language (i.e., associative similarity)”, which “suggests that GPT-3’s semantic network is organized around word meaning rather than their co-occurrence in text” Digutsch and Kosinski (2023).
@@
We augment existing research on the syntactic and semantic capabilities of LLMs, much of which is in terms of model performance (e.g., external manifestation rather than internal state), with under-the-hood examples of the sorts of things models can use tokens to encode. The approach developed here can be useful in many additional contexts and in future work.
@@
We are motivated to look into the representations of the most highly performant computational model of language to date [^44] LLMs in general, although we’re looking at a particular LLM, the RoBERTa MLM. because it provides a sketch of what might be minimally necessary to support human-like language performance. It also provides insight into a technology which is increasingly ubiquitous, which is important as we seek to understand what sorts of tasks these tools can be used in, and what sort of pitfalls they might afford. [^45] As research explores correspondences between human cognition and machine learning Binz et al. (2025), the horizon of possibilities for LLM applications in linguistics, cognitive science, and studies of human behavior continues to be re-evaluated. For this project, however, we focus strictly on human-like linguistic performance, not broader cognitive alignment. LLM linguistic abilities are sufficiently established for our purposes, despite their outputs not being perfectly human-like (if they were, then training on model-generated text would not cause collapse) Shumailov et al. (2024).
@@
We passed contiguous spans of text (a context window of 100 tokens on either side of the token of interest) sourced from a relatively small corpus of English-language fiction through a RoBERTa (large) Masked Language Model (MLM); we parsed the source text into tokens using a BPE tokenizer, the same approach RoBERTa uses. We then extracted the latent space data for our token of interest (in this paper, “ bank”, Fig. 9, “ run”, Fig. 40, and “ the”, Fig. 41, with leading spaces) by saving the embedding vector output by each layer of the model. We used consolidation strategies—UMAP and PCA—to boil down the large amount of data into something human-readable (through truncating and clustering). UMAP can result in “finer clustering than is necessarily present in the data,” which should be kept in mind when interpreting the resultant clusters McInnes, Healy, and Melville (2020, ). For more details of this process, see Sec. 25, Sec. 42.
@@
Figure 9: A condensed visualization of the content within instances of the “bank” token as it traveled through the RoBERTa MLM. Condensed representations of the trajectories themselves, in which the token vectors are 2D points moving from layer to layer within the model, are shown from various angles below the map figure. We condensed the trajectories using PCA, then UMAP, to create the map. The compass rose is purely aesthetic!
@@
Although the syntactic information often related to words which were sequentially very close, even adjacent, to the token within the utterance, this was not always the case. [^51] All of the information we could perceive is with respect to words pretty near the exemplar token, though, since our judgments were based on a 21 token context window (see Sec. 25). For example, the two segments of “the + [comparative adjective/adverb]” within the correlative comparative constructions525252e.g., think of phrases like, “the more, the merrier” in the were not necessarily very close together (see Fig. 41).
@@ 3.2 Objective functions
As one of the main forces shaping their cognition, [^74] For an example of the deep-rooted impact of objective function on cognition, Wolfe and Caliskan, 2022, found that “CLIP embeddings show that the high anisotropy observed by Ethayarajh (2019) is not the inevitable result of contextualization, nor even of a specific language modeling architecture, but is connected to the pretraining objective of the model” Wolfe and Caliskan (2022). (They mean pretraining as in before fine-tuning, so they still mean the objective function of the main intelligence.) The importance of the objective function to model cognition is also evident in results from transfer learning. exploring the space of objective functions is promising for allowing LLMs to get more out of the same data [^75] As in Zhuang et al., 2024, who propose that greater model efficiency could be achieved through more “cognitively aligned language learning technologies”, and found improvements in certain language representations when they modified the objective functions to include visual grounding Zhuang, Fedorenko, and Andreas (2024). and to solve some of the shortcomings that seem to come from such a stringent emphasis on producing plausible language above all else, a strategy that makes current LLMs silvertongued but ultimately unreliable and unhumanlike in terms of any behaviour outside those bounds (e.g., this could include aspects of Grice’s maxims, moral consistency, reciprocal self-disclosure, etc.) Grice (1989); Nunes et al. (2024). As one of the main forces shaping their cognition, [^76] For an example of the deep-rooted impact of objective function on cognition, Wolfe and Caliskan, 2022, found that “CLIP embeddings show that the high anisotropy observed by Ethayarajh (2019) is not the inevitable result of contextualization, nor even of a specific language modeling architecture, but is connected to the pretraining objective of the model” Wolfe and Caliskan (2022). (They mean pretraining as in before fine-tuning, so they still mean the objective function of the main intelligence.) The importance of the objective function to model cognition is also evident in results from transfer learning. exploring the space of objective functions is promising for allowing LLMs to get more out of the same data [^77] As in Zhuang et al., 2024, who propose that greater model efficiency could be achieved through more “cognitively aligned language learning technologies”, and found improvements in certain language representations when they modified the objective functions to include visual grounding Zhuang, Fedorenko, and Andreas (2024). and to solve some of the shortcomings that seem to come from such a stringent emphasis on producing plausible language above all else, a strategy that makes current LLMs silvertongued but ultimately unreliable and unhumanlike in terms of any behaviour outside those bounds (e.g., this could include aspects of Grice’s maxims, moral consistency, reciprocal self-disclosure, etc.) Grice (1989); Nunes et al. (2024).
@@ 6 Limitations
The main limitations are human fallibility, the many approximations we relied on (e.g., in creating lists of words, in using a small corpus of fiction, in using automated parsers, etc.) which were then composed on each other, and the many simplifying assumptions (moving between ‘orthography’ and ‘language’; considering broad families of architecture as basically similar enough, e.g., generative LLMs like the GPTs and MLMs like RoBERTa;128128128We would like to explore their differences in future work. In some ways, generating the next token and guessing a masked token are similar tasks, but surely there ought to be less flexibility for an MLM’s predictions. What is the impact of that constraint on the model’s gnogeography? As span increases, e.g., as the model is asked to supply one, two, n tokens, does its gnogeography approach an LLMs? How much does constraint vary with syntactic position and role? How much does constraint vary, for any model, with corpus? For example, when disorganized speech is plausible (as in poetry or certain psychiatric contexts) or across registers or topics, how does the model’s gnogeography change? Presumably topics or formats that are more free-wheeling ought to engender more uniform probability distributions. moving between English examples and abstract considerations of language as a technology). See Appendix 42 for further discussion of research limitations.

### v7

@@ 2.1 Exemplar tokenizations
A tokenization algorithm uses the training data it is given to determine what pieces (tokens) will be needed for future output. The popular tokenization design choice of maximizing orthographic efficiency allows as much data as possible to be processed by the LLM, presumably the motivation behind this strategy’s ubiquity. We trained one such tokenizer, a Byte-Pair Encoding (BPE) tokenizer Anthony Moi and Nicolas Patry (2023), with varying vocabulary sizes and training data to explore the details of the process, revealing some alignment between linguistic features and information processing by showing the human-like stages the tokens pass through. As a caveat, the training data chosen also determines the tokens; once vocabulary size becomes relatively large compared to the training data, tokens become too specific. The BPE implementation we used here was character-level, not byte-level, and we retained some (five) LLM-related special tokens as default items in the vocabulary. As a result, whatever our vocabulary size was fixed to, $N$, was in reality N5N-5, because it already included a few irrelevant (to us) tokens. We did not explicitly call any pre-tokenizers, in particular, the Whitespace pre-tokenizer, because that would mean the tokenizer would have automatically split on whitespace rather than treating the text agnostically as a continuous string of characters HuggingFace (2025). BPE is by no means the only tokenization algorithm used in LLMs, as aforementioned (Section 1.1), but our goal is to understand what kind of consequences follow for LLMs given that we have made tokenization – of any kind – an integral part of their architecture. Especially of interest to us are the implications of popular tokenization approaches that focus on parsing vast corpora of data efficiently, such as BPE, rather than on another goal, such as mimicking human language (which might make sense given the human-brain-inspiration of artificial neural nets and their application in LLMs to human language as primary data source) or on strategically designing cognition-scaffolding semantic primitives. [^28] The motivation underlying the creation of the Toki Pona conlang, and which would seem well-suited to neurosymbolic approaches.,  [^29] Future work could compare systematically across tokenization strategies.
@@
Figure 2: Tokens passing through word-like stages as the vocabulary size parameter increases, using a BPE tokenizer and the beginning of Alice in Wonderland. At the top of the figure is a syntax tree for our excerpt from Alice in Wonderland, and the colored paragraph next to it shows how the same excerpt could be tokenized by GPT-4o, underscoring the potential differences in representation between us and LLMs. Syntax tree via  Christodoulopoulos (2023); GPT-4o tokenization via  Duong (2023). The bar charts in the figure show the tokens used in the excerpt according to different tokenization schemas and vocabulary sizes. The height of the bars shows the count of each type of token in the excerpt. The charts on the left side of the figure show linguistically-informed tokenizations (that is, the excerpt if it were tokenized explicitly according to linguistic/ NLP categories, including words and subwords), and the GPT-4o tokenization, noting the similarities between it and the linguistically-informed tokenizations, especially the word-level tokenization. The charts on the right side of the figure show how increasing the vocabulary size causes the average token length to increase, but average count per token to decrease, as the tokens go from looking like individual letters, to subwords, to words. In the 700-token vocabulary, there are mostly words, but some phrases (‘she had’) and bound morphemes (‘ion’). The vocabulary size determines what the tokens will look like, and determines whether or not different linguistic types will be present in the vocabulary. For example, phrases and words will never occur in an extremely small vocabulary (given reasonably representative training data). The vocabulary then determines how the model represents all text as vectors.
The tokenizer’s vocabulary size hyperparameter sets the total number of tokens that will be known to the model. Due to the way the BPE tokenization algorithm works, as the vocabulary size increases, the maximum length of the tokens also increases (an initial base vocabulary covering the fundamental units, usually characters or bytes, is expanded through a series of pairwise merges). Therefore, in iteratively increasing the vocabulary size, we must313131in the logic of the Intermediate Value Theorem at some point pass through token lengths that are the lengths of human-meaningful units, e.g., word length, but the tokens do not need to actually be those units, e.g., words.
@@
Figure 3: The tokenization (Y-axis) of the first sentence from Alice in Wonderland – “Alice was beginning to get very tired of sitting by her sister on the bank, and of having nothing to do: once or twice she had peeped into the book her sister was reading, but it had no pictures or conversations in it, ‘and what is the use of a book,’ thought Alice ‘without pictures or conversations?”’ – by a BPE tokenizer trained on the wikitext dataset and that sentence, as the vocabulary size (X-axis) increases from 2,000 to 30,000, then additional much larger vocabularies (the figure can be explored in a PDF viewer). The tokens at vocabulary size of 1,000 were individual characters; there were around 3 times as many tokens in the sentence as at vocabulary size of 2,000, so imagine a tower on the extreme left that is 3 times as tall as the one shown. Likewise, on the extreme right would eventually be the sentence as a single token. We skipped some vocabulary sizes along the x-axis: it is not evenly spaced (intermediate behaviour between the largest vocabulary sizes was omitted for legibility, but can be reasonably extrapolated by the reader). Middle English translation via OpenL . The short tokens, while combinatorily potent and requiring the model to memorize only a small vocabulary, engender prolific output: the ratio of total tokens per generated text are essentially as high as possible given the source material. On the other hand, the large tokens, while efficient in terms of total tokens per generated text (in the limit, 1:1), are far too specific (the model would need to memorize an essentially infinite vocabulary in order to output general purpose text). In other words, compared to the original sentence, the different tokenizations show how it gets chopped up as we increase the vocabulary size, revealing a tradeoff between tokens as individual characters, in which case, we’d need to encode all meaning in context alone, versus tokens as full sentences without generalizable meaning, in which case, we’d need an infinitely large vocabulary to produce sensible output.
@@ 2.2 Exemplar vocabularies
In this section, we examine a set of LLM (including both autoregressive and masked language models) vocabularies to explore what tokens look like ‘in the wild’.
GPT-4o is a recent, prominent, proprietary flagship model OpenAI et al. (2024); Shahriar et al. (2024), and we take its vocabulary to be reasonably representative of what a good vocabulary currently looks like. We can see in Figure 2 that GPT-4o uses tokens that look similar to the word-like stage (achieved through linguistic-agnostic tokenizing) and like tokens generated based on linguistic structures (morphemes, lemmas, words), consistent with our expectation.
However, relatively little information is available on what tokens look like more generally, such as what tokens look like in vocabularies across various in-use models. An exception is (the very fun) work of Zhemchuzhina et al., 2022, which delves into what kinds of tokens occur in a vocabulary, and comes up with the following categories: (1) atom, the smallest discrete element, which is probably coming from a closed set (as in the letters of an alphabet in written language); (2) pragma, which consists of atoms, represents part of an idea, and has multiple meanings (which seemingly would correspond to morphemes); (3) idea, a sequence of pragmas whose meaning is a single specific concept (which they relate to utterances and phrases). Due to the vocabulary size of most models, they point out that most tokens correspond to atoms and pragmas rather than ideas Zhemchuzhina, Filippov, and Yamshchikov (2022).
Relatedly, we expect that function words (e.g., “from”, “this”), generally short, frequent, and considered a fairly small closed class in English Brysbaert et al. (2016), are therefore more likely to be tokenized (i.e., even occur in small vocabularies) compared to the majority of content words. [^35] This seems to be true as shown in Figure 8.
We extend this work by examining a group of in-use LLM vocabularies (including models with both autoregressive and masked training objectives) to see what their tokens actually look like. We specifically seek commonalities between human language and tokens, since we suspect emerging best practices should be nudging vocabularies in the human-like direction, but our primary goal is to gain insight into what the tokens in real, reasonably widely-used models look like.
@@ 2.3 Extispicy: exemplar tokens
We augment existing research on the syntactic and semantic capabilities of LLMs, much of which is in terms of model performance (e.g., external manifestation rather than internal state), with under-the-hood examples of the sorts of things models can use tokens to encode. Previous work found semantic information was a dominant force within the GPT-3 LLM’s internal organization of language. Digutsch and Kosinski, 2023, found that “GPT-3’s semantic activation is better predicted by similarity in words’ meaning (i.e., semantic similarity) rather than their co-occurrence in the language (i.e., associative similarity)”, which “suggests that GPT-3’s semantic network is organized around word meaning rather than their co-occurrence in text” Digutsch and Kosinski (2023). We add to this result with detailed illustrations of some of the structures supporting this internal organization of language. The approach developed here can be useful in many additional contexts and in future work.
@@
We are motivated to look into the representations of the most highly performant computational model of language to date [^46] By which we mean, LLMs in general, although we’re looking at a particular LLM, the RoBERTa MLM. because it provides a sketch of what might be minimally necessary to support human-like language performance. It also provides insight into a technology which is increasingly ubiquitous, which is important as we seek to understand what sorts of tasks these tools can be used in, and what sort of pitfalls they might afford. [^47] As research explores correspondences between human cognition and machine learning Binz et al. (2025), the horizon of possibilities for LLM applications in linguistics, cognitive science, and studies of human behavior continues to be re-evaluated. For this project, however, we focus strictly on human-like linguistic performance, not broader cognitive alignment. LLM linguistic abilities are sufficiently established for our purposes, despite their outputs not being perfectly human-like (if they were, then training on model-generated text would not cause collapse) Shumailov et al. (2024).
@@
We passed contiguous spans of text (a context window of 100 tokens on either side of the token of interest) sourced from a relatively small corpus of English-language fiction through a RoBERTa (large) Masked Language Model (MLM); we parsed the source text into tokens using a BPE tokenizer, the same approach RoBERTa uses. We then extracted the latent space data for our token of interest (in this paper, “ bank”, Fig. 9, “ run”, Fig. 40, and “ the”, Fig. 41, with leading spaces) by saving the embedding vector output by each layer of the model. We used consolidation strategies—UMAP and PCA—to boil down the large amount of data into something human-readable (through truncating and clustering). UMAP can result in “finer clustering than is necessarily present in the data,” which should be kept in mind when interpreting the resultant clusters McInnes, Healy, and Melville (2020, ). For more details of this process, see Sec. 25, Sec. 42. Our reasons for using the RoBERTa model are discussed in the Appendix (Sec. 7). These particular gnogeographic maps are unique to RoBERTa and the text we chose to use; other models and other contexts would both yield different maps. What we are exploring is what these internal organizations can look like, and what kind of strategies a model can use to organize. Although the model’s architecture (including its training objective) and how it was trained will each impact the resultant gnogeographic maps, given that LLMs as a group exhibit impressive natural language competency compared to older models and do have significant commonalities between them, we don’t see any reason to think these gnogeographic maps would be irrelevant outside of this specific combination of model and corpus.
@@
Figure 9: A condensed visualization of the content within instances of the “ bank” token as it traveled through the RoBERTa MLM. Condensed representations of the trajectories themselves, in which the token vectors are 2D points moving from layer to layer within the model, are shown from various angles below the map figure. We condensed the trajectories using PCA, then UMAP, to create the map. We took the position of the “ bank” token vectors after each layer of the model, then condensed those trajectories with PCA, then used UMAP to cluster the condensed trajectories. The maps show the clusters according to the first two UMAP dimensions. We sampled sentences at points within the clusters and annotated the map with our interpretation of what was in that cluster. The end result reveals that the clusters distinguish the financial sense of ‘bank’ from ‘databank’ or ‘river bank’, and that the top-level meanings (financial, river) can be further subdivided, e.g. the financial meaning is divided into the physical institution of a bank versus the abstract financial entity and associated functionality. More details on the contents of the gnogeographic maps are provided in the main body of the paper, including discussion of embodiment, time, space, Aktionsart, prototypical usage, syntax, formatting, and some very unhuman-like potential ‘meanings’ such as a vehicle sense for “ the”! The compass rose, sea monster, and ship are purely aesthetic.
@@
Although the syntactic information often related to words which were sequentially very close, even adjacent, to the token within the utterance, this was not always the case. [^53] All of the information we could perceive is with respect to words pretty near the exemplar token, though, since our judgments were based on a 21 token context window (see Sec. 25). For example, the two segments of “the + [comparative adjective/adverb]” within the correlative comparative constructions545454e.g., think of phrases like, “the more, the merrier” in the were not necessarily very close together (see Fig. 41).
@@ 3.2 Objective functions
As one of the main forces shaping their cognition, [^76] For an example of the deep-rooted impact of objective function on cognition, Wolfe and Caliskan, 2022, found that “CLIP embeddings show that the high anisotropy observed by Ethayarajh (2019) is not the inevitable result of contextualization, nor even of a specific language modeling architecture, but is connected to the pretraining objective of the model” Wolfe and Caliskan (2022). (They mean pretraining as in before fine-tuning, so they still mean the objective function of the main intelligence.) The importance of the objective function to model cognition is also evident in results from transfer learning. exploring the space of objective functions is promising for allowing LLMs to get more out of the same data [^77] As in Zhuang et al., 2024, who propose that greater model efficiency could be achieved through more “cognitively aligned language learning technologies”, and found improvements in certain language representations when they modified the objective functions to include visual grounding Zhuang, Fedorenko, and Andreas (2024). and to solve some of the shortcomings that seem to come from such a stringent emphasis on producing plausible language above all else, a strategy that makes current LLMs silvertongued but ultimately unreliable and unhumanlike in terms of any behaviour outside those bounds (e.g., this could include aspects of Grice’s maxims, moral consistency, reciprocal self-disclosure, etc.) Grice (1989); Nunes et al. (2024).
@@ 6 Limitations
The main limitations are human fallibility and their downstream functional implications (limited resources, limited concentration, limited knowledge, limited experience), the many approximations we relied on (e.g., in creating lists of words, in using a small corpus of fiction, in using automated parsers, etc.) which were then composed on each other, and the many simplifying assumptions (moving between ‘orthography’ and ‘language’; considering broad families of architecture as basically similar enough, e.g., generative LLMs like the GPTs and MLMs like RoBERTa;128128128We would like to explore their differences in future work. In some ways, generating the next token and guessing a masked token are similar tasks, but surely there ought to be less flexibility for an MLM’s predictions. What is the impact of that constraint on the model’s gnogeography? As span increases, e.g., as the model is asked to supply one, two, n tokens, does its gnogeography approach an LLMs? How much does constraint vary with syntactic position and role? How much does constraint vary, for any model, with corpus? For example, when disorganized speech is plausible (as in poetry or certain psychiatric contexts) or across registers or topics, how does the model’s gnogeography change? Presumably topics or formats that are more free-wheeling ought to engender more uniform probability distributions. moving between English examples and abstract considerations of language as a technology; using terms like ‘word-like’, ‘cognition’, and ‘meaning’ without providing precise definitions). We are happy to discuss why we made the choices we did with interested parties. See Appendix 42 for further discussion of research limitations.